# Diversity as a Reward: Fine-Tuning LLMs on a Mixture of Domain-Undetermined Data

### v1

@@ 3.3 Experimental Observations
Table 1 presents comprehensive evaluations across seven benchmarks, where the notation Inter-Diversity (X-Y) indicates samples ranked in the top (100-Y)to (100-X)of cross-domain similarity scores. Due to space constraints, we present only the results for the top 20%, middle 20%, and bottom 20%. More results can be found in Appendix D.7.
Diversity-Aware Performance: Our diversity-controlled selections reveal two critical observations:
  • Varied Improvement Patterns: Both models demonstrate marked improvements over Raw distributions across all diversity conditions, but the effects of their improvements vary. For Llama3.1-8B, Inter-D (80-100) achieves 38.98 average accuracy (+3.12 over Raw), outperforming the Random baseline by 1.71, while Inter-D (0-20) is below Random of 0.09.
@@
Despite existing positive improvements on overall performance, the optimal distribution parameters exhibit model-dependent variability. This parameter sensitivity suggests the existence of multiple local optima in the diversity-performance landscape. Two constraints merit consideration for real-world applications. (1) Label Dependency: The studied heuristic strategies Inter-Diversity and Intra-Diversity currently require domain-labeled data for centroid calculation. (2) Distribution Transiency: The optimal diversity parameters (e.g., 80-100 vs. 40-60) show sensitivity across tasks and models, necessitating automated and potentially costly configuration search.
Table 1: Performance of Llama3.1-8B and Qwen2-7B on various downstream task benchmarks under different constructed Inter-Diversity (Inter-D) and Intra-Diversity (Intra-D) distributions.
@@
Random 21.99 64.83 74.72 55.70 14.50 5.10 24.09 37.27
Inter-D (0-20) 19.28 65.79 74.44 54.90 6.50 4.30 35.06 37.18
Inter-D (40-60) 23.70 65.14 74.86 56.40 17.15 5.00 24.40 38.09
Inter-D (80-100) 23.76 64.43 75.20 56.40 15.05 4.50 33.54 38.98
Intra-D (0-20) 22.08 65.08 75.00 54.70 16.20 4.40 33.54 38.71
Intra-D (40-60) 22.12 64.74 74.87 54.00 16.00 6.00 27.44 37.88
Intra-D (80-100) 19.78 64.77 74.51 56.50 13.00 5.20 37.50 38.75
@@
Random 13.28 58.27 73.00 75.35 35.36 52.20 63.72 53.02
Inter-D (0-20) 15.18 59.28 73.34 74.50 34.94 53.10 68.60 54.13
Inter-D (40-60) 14.62 58.58 73.35 72.50 34.50 52.50 61.28 52.47
Inter-D (80-100) 9.30 57.72 73.14 74.60 28.00 51.30 63.42 51.07
Intra-D (0-20) 12.64 58.54 73.35 75.10 8.75 51.10 61.59 48.72
Intra-D (40-60) 15.24 58.57 73.12 74.70 32.50 51.80 64.02 52.85
Intra-D (80-100) 11.91 57.88 73.29 75.00 36.05 52.50 66.16 53.25
4 Theoretical Analysis and Insights
In this section, we first formalize the optimization dynamics of multi-domain data aggregation for enhancing LLMs’ comprehensive capabilities. Building on the empirical observations from Sec. 3, we then establish theoretical connections between distributional diversity and emergent model behaviors. Finally, we demonstrate the fundamental limitation of explicit diversity optimization in undetermined-domain scenarios.
4.1 Ideal Optimization Formulation
During the training process of LLMs, various sources of labeled data are often collected to enhance models’ capabilities across multiple dimensions. Let $\mathcal{K}=\{1,\ldots,K\}$ represent the set of labeled domains, where $K$ is determined by real-world task requirements (e.g., $K=4$ in our experiments). For each domain $k\in\mathcal{K}$, data samples $\mathbf{x}_{i}^{(k)}\in\mathcal{D}_{k}$ are generated according to a distribution $\mathcal{D}_{k}$. The data distributions are assumed mutually distinct across different domains, leading to an ideal I.I.D. per-domain hypothesis: each domain $k$ corresponds to an independent model $h_{k}\in\mathcal{H}$. The standard optimization goal thus can be:
$\small\forall k\in\mathcal{K},\quad\min_{h_{k}\in\mathcal{H}}\mathcal{L}_{\mathcal{D}_{k}}(h_{k}),$ (6)
where $\mathcal{L}_{\mathcal{D}_{k}}(h_{k})=\mathbb{E}_{(x,y)\sim\mathcal{D}_{k}}\left[l(h_{k}(x),y)\right]$ is the empirical risk given a specific loss function $l$.
Counterintuitive Findings
If the distributions $\{\mathcal{D}_{k}\}_{k=1}^{K}$ are strictly I.I.D., the risk $\mathcal{L}_{\mathcal{D}_{k}}(h_{k})$ should depend only on $\mathcal{D}_{k}$. However, numerous previous studies (Zhao et al., 2024a; Tirumala et al., 2023) have demonstrated the presence of synergistic and antagonistic effects between different datasets. This is further demonstrated in Sec. 3 (Fig. 1) that certain domain mixtures lead to catastrophic performance degradation (e.g. Intra-D (0-20) in MATH with Qwen2). This contradicts the above I.I.D. hypothesis, suggesting that the simplified and conventional formulation in Eq (6) fails to capture cross-domain interference.
4.2 Introducing Mixture of Underlying Distributions
Inspired by the counterintuitive findings, we posit that each domain distribution $\mathcal{D}_{k}$ arises from the mixture of $M$ latent distributions $\{\tilde{\mathcal{D}}_{m}\}_{m=1}^{M}$ representing foundational LLM capabilities ($M\ll K$ in practice). This leads to our core statistical assumption:
Assumption 4.1 (Latent Capability Structure).
For each observed domain $k\in\mathcal{K}$, its data distribution decomposes into $M$ latent capability distributions:
$\small\mathcal{D}_{k}=\sum_{m=1}^{M}\pi_{km}\tilde{\mathcal{D}}_{m},\quad\sum_{m=1}^{M}\pi_{km}=1,$ (7)
where $\tilde{\mathcal{D}}_{m}$ indicates the $m$-th foundational capability. The weights $\pi_{km}$ reflect domain-specific capaility composition.
To analyze how data optimization interacts with latent capability distributions, we further posit that each foundational capability admits an optimal configuration:
Assumption 4.2 (Capability Optimality).
Each foundational capability admits a unique optimal predictor:
$\small\exists h_{\theta_{m}^{*}}\in\mathcal{H}\text{ s.t. }\theta_{m}^{*}=\mathop{\mathrm{argmin}}_{\theta}\mathbb{E}_{(x,y)\sim\tilde{\mathcal{D}}_{m}}[l(h_{\theta}(x),y)].$ (8)
The loss $l$ is strongly convex, holding for cross-entropy or MSE, where $h_{\theta_{m}}$, termed the component model, represents a model’s manifestation of specific foundational capabilities.
This leads to our key result formalizing how real-world training data activates latent capabilities via component models:
Proposition 4.3.
Under Assumption 4.1-4.2, the minimizer of the cross-domain average risk
$\small\min_{\{\theta_{m}\}}\mathbb{E}_{k\sim\mathcal{K}}\left[\mathcal{L}_{\mathcal{D}_{k}}\left(\textstyle\sum_{m=1}^{M}\pi_{km}h_{\theta_{m}}\right)\right]$ (9)
admits a solution $\{\theta_{m}^{*}\}_{m=1}^{M}$ where each $h_{\theta_{m}^{*}}$ optimizes the corresponding foundational capability $\tilde{\mathcal{D}}_{m}$.
Remark Proposition 4.3 reveals an underlying fact: optimal or suboptimal solutions can be achieved through data selection to enhance the overall model performance. This is further validated by the role of the coefficients $\pi_{km}$, which act as capability selectors to configure domain-specific behavior, demonstrating the impact of data composition.
4.3 Diversity Influence on the Optimization
Building upon Proposition 4.3, we analyze the intrinsic relationship between diversity and capability composition. The mixture coefficients $\pi_{km}$ inherently govern both inter- and intra-domain diversity formulated in Sec. 3.2, differing in their geometric properties in the latent capability space.
Centroids with Coefficients Let $\mathcal{C}_{m}\in\mathbb{R}^{d}$ denote the centroid vector of latent capability $\tilde{\mathcal{D}}_{m}$ in the embedding space. The domain centroid $\mathcal{C}_{k}$ can be expressed as:
$\small\mathcal{C}_{k}=\sum_{m=1}^{M}\pi_{km}\mathcal{C}_{m}.$ (10)
This linear combination implies that the inter- and intra-diversity are determined by $\pi_{km}$ configurations:
Proposition 4.4 (Diversity Decomposition).
The inter-diversity metric decomposes into:
$\small\Phi_{\text{inter}}=\sum_{m=1}^{M}\sum_{n=1}^{M}\lambda_{m}\lambda_{n}\|\mathcal{C}_{m}-\mathcal{C}_{n}\|_{2},$ (11)
where $\lambda_{m}=\mathbb{E}_{k}[\pi_{km}]$. And the intra-diversity satisfies:
$\small\Phi_{\text{intra}}^{(k)}\leq\sum_{m=1}^{M}\pi_{km}\|\mathcal{C}_{m}-\mathcal{C}_{k}\|_{2}^{2}+\mathbb{E}_{m\sim\pi_{k}}[\text{Var}(\tilde{\mathcal{D}}_{m})].$ (12)
The inter-diversity result follows from substituting $\mathcal{C}_{k}=\sum\pi_{km}\mathcal{C}_{m}$ into $\mathbb{E}_{k\neq l}\|\mathcal{C}_{k}-\mathcal{C}_{l}\|$ and applying Jensen’s inequality. The intra-diversity bound combines the variance decomposition within each latent capability. Full proof is provided in Appendix C.2.
4.4 Theoretical Insights
Collectively, results in this section provide a unified framework to explain the varied patterns and existence of peak performance observed in Sec. 3.3, providing theoretical insights and guidance for designing our following algorithm dealing with undetermined data.
Diversity’s Influence on LLMs’ Capability: Based on Proposition 4.4, we conjecture that the Inter-/Intra-Diversity, which can be explicitly modeled by $\pi$, also governs $\pi$-variation control for component model mixing and thus the final overall performance. Proposition 4.3 further suggests that improper diversity levels may induce model suboptimality by constraining overall capability, consistent with the distribution-dependent empirical results in Sec. 3.
Robust Diversity Selection Strategies: Based on Proposition 4.4, it can be observed that averaging the $\pi_{km}$ helps to mitigate the missing of underlying optimal predictor $\theta_{m}^{*}$. From this view, the random baseline can gain relatively stable improvements in expectation. A more robust yet simple strategy can be derived: ensemble across candidate models trained on multiple data pools with distribution-varied inter- and intra-diversity, e.g., via techniques like voting or model weights averaging. This can be corroborated by results in Table 1, such as the case for Llama3.1-8B.
Challenges in Undetermined-Domain Scenarios: However, all the selection strategies discussed so far face intractable solutions in undetermined domains with explicit domain labels, as Proposition 4.4 technically revealed: (1) ambiguous domain boundaries, e.g., blended instruction types; (2) partial overlaps in latent capabilities; and (3) absence of centroid priors $\{\mathcal{C}_{m}\}$. These intrinsic constraints motivate our model-aware diversity reward in the following Sec. 5, which harnesses domain-diversity awareness into reward modeling without requiring explicit parameterization for such structural assumptions.
5 DaaR: Diversity as a Reward
To address the challenges identified and leverage the insights gained in previous Sec. 3 and Sec. 4, we establish a data selection method DaaR guided by diversity-aware reward signals. It comprises three key components illustrated in subsequent sections: (1) model-aware centroid synthesis, (2) two-stage training with reward probe, and (3) diversity-driven data selection.
5.1 Model-Aware Training Data
Model-Aware Centroid Construction
  • Phase 1 - Seed Generation: For each domain $k$, generate 5 seed samples $\mathcal{S}_{k}^{(0)}$ via zero-shot prompting with domain-specific description templates, establishing initial semantic anchors. The prompt details and ablation of choices on sample number are provided in Appendix B.7.
  • Phase 2 - Diversity Augmentation: Iteratively expand $S_{k}^{(t)}$ through context-aware generation, conditioned on a sliding window buffer with 3 random anchors sampled from the $(t-1)$ iteration $S_{k}^{(t-1)}$. The generated sample $\mathbf{x^{\prime}}$ will be admitted during the iteration via geometric rejection sampling:
    $\max_{\mathbf{x}\in S_{k}^{(t-1)}}\cos(\mathcal{M}_{\text{ebd}}(\mathbf{x^{\prime}}),\mathcal{M}_{\text{ebd}}(\mathbf{x}))<\tau=0.85,$ (13)
    where $\mathcal{M}_{\text{ebd}}(\cdot)$ indicates the embedding layer outputs of the given LLM. This process terminates when $|S_{k}|=30$. More analysis and ablation on these choices are provided in the Appendix. In Appendix D.3, we found that the adopted number of samples sufficiently leads to a stable convergence and larger data quantity does not impact the final centroids. In Appendix D.4, we found that this synthetic data has the ability to produce domain-representative data with clear distinction. Interestingly, the generated content consistently exhibits the greatest divergence between common sense, reasoning, and coding domains across model architectures and parameters.
The domain centroid is then computed from the final augmented set $\mathcal{S}_{k}$ using the model’s knowledge prior:
$\mathcal{C}_{k}=\frac{1}{|\mathcal{S}_{k}|}\sum_{\mathbf{x}_{i}\in\mathcal{S}_{k}}\mathcal{M}_{\text{ebd}}(\mathbf{x}_{i}).$ (14)
This captures the LLM’s intrinsic feature space geometry while eliminating dependency on human annotations.
Domain-Aware Clustering
We then automatically construct pseudo-labels for the given data samples based on the previously synthesized centroids $\{C_{k}\}_{k=1}^{K}$. Take them as fixed prototypes, we perform constrained k-means clustering in the embedding space:
$\arg\min_{\{S_{k}\}}\sum_{k=1}^{K}\sum_{\tilde{x}\in S_{k}}\|\mathcal{M}_{\text{ebd}}(\tilde{x})-\mathcal{C}_{k}\|^{2}.$ (15)
This produces the seed dataset $\mathcal{D}_{\text{probe}}$ containing pseudo-labels $\{\tilde{y}_{i}\}_{i=1}$ where $\tilde{y}_{i}\in\{1,\ldots,K\}$, with model-induced and embedding-derived domain label assignments.
5.2 Training for Self-Rewarding Abilities
Stage 1: Domain Discrimination
The proposed DaaR then establishes model-aware domain discrimination abilities through a multi-layer perceptron (MLP) probe module, $\psi_{\text{dom}}$, attached to the layer-3 of the LLMs $\mathcal{M}_{3}(\tilde{x})$. The probe will be trained meanwhile all the parameters of the LLM are frozen. This achieves a preferable balance between effectiveness and cost, with detailed analysis regarding the choice of Layer-3 presented in Appendix D.5. Specifically, with pseudo-label $\tilde{y}$, we can compute domain probabilities as:
$p_{k}(\tilde{x})=\text{softmax}\left(\psi_{\text{dom}}(\mathcal{M}_{3}(\tilde{x}))\right),\quad\psi_{\text{dom}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{K},$ (16)
where $\psi_{\text{dom}}$ is optimized via cross-entropy loss function:
$\mathcal{L}_{\text{dom}}=-\frac{1}{|\mathcal{D}_{\text{probe}}|}\sum_{(\tilde{x},\tilde{y})\in\mathcal{D}_{\text{probe}}}\sum_{k=1}^{K}\mathbb{I}_{[k=\tilde{y}]}\log p_{k}(\tilde{x}),$ (17)
where $\mathbb{I}_{[k=\tilde{y}]}$ denotes the indicator function. We employ single-sample batches with the AdamW optimizer to prevent gradient averaging across domains. Training consistently converges and achieves 92.7% validation accuracy on domain classification, as shown in Fig. 2 (a).
Figure 2: Training loss and validation process of the two stages of DaaR on Qwen2-7B, detailed results are in Appendix D.6.
Stage 2: Diversity Rewarding
Building on the stabilized domain probe module, we quantify sample-level diversity through predictive entropy:
$H(\tilde{x})=-\sum_{k=1}^{K}p_{k}(\tilde{x})\log p_{k}(\tilde{x}).$ (18)
To enable efficient reward computation during data selection, we then train another 5-layer MLP $\psi_{\text{div}}$ to directly estimate $H(\tilde{x})$ from $\mathcal{M}_{3}(\tilde{x})$:
$\hat{H}(\tilde{x})=\psi_{\text{div}}(\mathcal{M}_{3}(\tilde{x})),\quad\psi_{\text{div}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}.$ (19)
This diversity probe module $\psi_{\text{div}}$ shares $\psi_{\text{dom}}$’s architecture up to its final layer (replaced with single-output regression head), trained using entropy-scaled MSE:
$\mathcal{L}_{\text{div}}=\frac{1}{|\mathcal{D}_{\text{probe}}|}\sum_{\tilde{x}\in\mathcal{D}_{\text{probe}}}\left(\hat{H}(\tilde{x})-H(\tilde{x})\right)^{2}.$ (20)
This module is also well-converged as shown in Fig. 2 (b).
Table 2: Evaluation results of Llama3.1-8B, Qwen2-7B, and Qwen2.5-7B across various downstream task benchmarks. DaaR demonstrates superiority in average performance (AVG) compared to other baselines.
Models Distribution Common Sense Reasoning Mathematics Coding Avg
NQ TriviaQA Hellaswag GSM8K MATH MBPP HumanEval
Llama3.1-8B Raw 14.13 65.90 74.62 54.80 7.90 5.00 28.66 35.86
Random 21.99 64.83 74.72 55.70 14.50 5.10 24.09 37.27
Instruction Len 15.34 63.60 73.73 54.00 15.40 3.60 30.80 36.64
Alpagasus (Chen et al., 2024d) 21.57 64.37 74.87 55.20 17.65 4.60 16.16 36.34
Instag-Best (Lu et al., 2024) 18.12 64.96 74.01 55.70 15.50 4.80 37.81 38.70
SuperFilter (Li et al., 2024a) 22.95 64.99 76.39 57.60 6.05 2.60 40.55 38.73
Deita-Best (Liu et al., 2024b) 15.58 64.97 74.21 55.00 13.05 4.60 34.46 37.41
DaaR (Ours) 20.08 64.55 74.88 54.8 15.30 4.70 37.50 38.83
Qwen2-7B Raw 8.03 59.58 73.00 78.00 5.70 5.00 60.98 41.47
Random 13.28 58.27 73.00 75.35 35.36 52.20 63.72 53.02
@@ 5.2 Training for Self-Rewarding Abilities
Alpagasus (Chen et al., 2024d) 13.67 57.94 73.04 73.90 32.30 51.40 63.41 52.24
Instag-Best (Lu et al., 2024) 9.51 58.50 73.06 74.70 35.35 51.90 64.70 52.53
SuperFilter (Li et al., 2024a) 19.16 58.98 72.99 73.70 30.10 52.40 58.85 52.31
Deita-Best (Liu et al., 2024b) 16.41 57.80 72.70 76.10 29.05 52.40 64.63 52.73
@@
Alpagasus (Chen et al., 2024d) 10.40 57.87 72.92 77.20 18.75 61.80 65.55 52.07
Instag-Best (Lu et al., 2024) 11.08 58.40 72.79 76.40 16.40 62.90 70.43 52.63
SuperFilter (Li et al., 2024a) 13.54 58.51 72.89 79.30 11.35 39.50 65.25 48.62
Deita-Best (Liu et al., 2024b) 10.50 58.17 73.14 74.60 16.60 62.00 72.26 52.47

### v2

@@ 3.3 Experimental Observations
Table 1 presents comprehensive evaluations across seven benchmarks, where the notation Inter-Diversity (X-Y) indicates samples ranked in the top (100-Y)to (100-X)of cross-domain similarity scores. Due to space constraints, we present only the results for the top 20%, middle 20%, and bottom 20%. Detailed performance results on various downstream tasks are shown in Appendix H.4. Our diversity-controlled selections reveal two observations:
Table 1: Performance of Llama3.1-8B and Qwen2-7B on Avg with different constructed Inter-Diversity and Intra-Diversity distributions.
Models Distribution Avg
Llama3.1 Raw 35.86
Full (40K) 37.22
Random (8K) 37.27
Inter-D (0-20) 37.18
Inter-D (40-60) 38.09
Inter-D (80-100) 38.98
Intra-D (0-20) 38.71
Intra-D (40-60) 37.88
Intra-D (80-100) 38.75
Qwen2 Raw 41.47
Full (40K) 53.01
Random (8K) 53.02
Inter-D (0-20) 54.13
Inter-D (40-60) 52.47
Inter-D (80-100) 51.07
Intra-D (0-20) 48.72
Intra-D (40-60) 52.85
Intra-D (80-100) 53.25
  • Varied Improvement Patterns: Both models demonstrate marked improvements over Raw distributions across diversity conditions, but the effects of their improvements vary. For Llama3.1-8B, Inter-D (80-100) achieves 38.98 average accuracy (+3.12 over Raw), outperforming the Random baseline by 1.71, while Inter-D (0-20) is below Random.
@@
Remarks Despite existing positive improvements on overall performance, two constraints merit consideration for real-world applications. (1) Distribution Transiency: The optimal diversity parameters (e.g., 80-100 vs. 40-60) show sensitivity across tasks and models, necessitating automated and potentially costly methods. (2) Label Dependency: The studied heuristic strategies Inter-Diversity and Intra-Diversity currently require domain-labeled data for centroid calculation.
3.4 Theoretical Perspective
To explain why diversity serves as a crucial factor in LLM’s cross-domain performance and addresses distribution transiency, we present a theoretical analysis based on important sampling. We first formulate the data selection dynamics and then incorporate diversity into it, thereby revealing the mechanistic basis and offering insights for our method.
Importance Sampling
We assume that there exists an ideal target distribution $q$ that minimizes the multi-domain loss in Eq. (1). Hence, the process of data selection can be framed as aligning the source distribution $p$ (empirical distribution of $\mathcal{D}$) toward the optimal distribution $q$. To bridge this distribution shift, we adopt an importance sampling perspective xie2023data ; xie2023doremi ; xialess where the optimal subset corresponds to reweighting samples by the density ratio $\frac{q(x)}{p(x)}$:
[…]
The domains of LLM’s data are commonly defined by textual descriptions, which inherently constrain the semantic scope. Thus we make the following assumption regarding distributions $p$ and $q$:
Assumption 3.1 (Shared Support).
For any given domain index $c\in\{1,\dots,K\}$, the support of $p$ and $q$ is shared, formalized as: $p(x|c=k)=q(x|c=k),\quad k=1,\dots,K.$
Assumption 3.1 implies that the discrepancy between the source distribution $p$ and the target distribution $q$ arises solely from differences in domain proportions, leading to the following proposition.
Proposition 3.2.
The diversity is entirely attributable to the relative weights assigned to each domain:
$\frac{q(x)}{p(x)}=\sum_{k=1}^{K}\lambda_{k}p(c{=}k|x),\lambda_{k}:=\frac{q(c=k)}{p(c=k)}$ (5)
The derivation is detailed in Appendix E. This decomposition reveals that the Inter-Diversity and Intra-Diversity are governed by $\lambda_{k}$ and $p(c{=}k|x)$, leading to the following analysis.
Why Overlook Diversity is Suboptimal
We define $k^{*}(x):=\arg\max_{k}p(c=k|x)$ as the prediction of the classifier $p(c=k|x)$. Consider a special case where the classifier yields deterministic predictions, i.e. $p(c=k^{*}(x)|x)=1,\forall x\in\mathcal{D}$. In this case, the importance weight reduces to $\frac{q(x)}{p(x)}=\sum_{k=1}^{K}\lambda_{k}p(c=k|x)=\lambda_{k^{*}(x)}$. Therefore, importance sampling can be applied within:
$\small\sum_{x\in\mathcal{D}}\frac{q(x)}{p(x)}\ell(x;\theta)=\sum_{x\in\mathcal{D}}\lambda_{k^{*}(x)}\ell(x;\theta)=\sum_{k=1}^{K}\ \sum_{x\in\mathcal{D}:k^{*}(x)=k}\lambda_{k*(x)}\ell(x;\theta).$ (6)
This formulation is commonly used when domain labels are available, under the implicit assumption that domains are mutually exclusive. However, such an assumption rarely holds in the context of LLM training, as cross-domain data may either synergize or conflict zhao2024beyond ; xie2023doremi (evidence in Table 1). Hence, neglecting diversity in data selection potentially leads to suboptimal overall performance.
Guidance to Our Method
Under the domain-undetermined scenario, our proposed method DaaR operationalizes the gained insights by (1) constructing pseudo-labels (Section 4.1) align with LLM-aware generated seed, (2) explicitly modeling the classifier $p(c{=}k|x)$ through a domain discrimination probe (Section 4.2), and (3) using the probe’s predictive entropy (Section 4.2) as a proxy for diversity, estimating the ability of classifier $p(c{=}k|x)$ to select data toward the target distribution $q$.
4 DaaR: Diversity as a Reward
To address the challenges identified in Sec. 3.3 and leverage the insights gained in Sec. 3.4, we establish a data selection method DaaR guided by diversity-aware reward signals. It comprises three key components illustrated in Fig. 1: (1) model-aware centroid synthesis, (2) two-stage training with reward probe, and (3) diversity-driven data selection.
4.1 Model-Aware Training Data
Model-aware Centroid Construction
  • Phase 1 - Seed Generation: For each domain $k$, generate seed samples $\mathcal{S}_{k}^{(0)}$ via zero-shot prompting with domain-specific description templates generated by LLM itself, alongside with minor injection from downstream task samples, establishing initial semantic anchors.
  • Phase 2 - Diversity Augmentation: Iteratively expand $S_{k}^{(t)}$ through context-aware generation, conditioned on a sliding window buffer with 3 random anchors sampled from the $(t-1)$ iteration $S_{k}^{(t-1)}$. The generated sample $\mathbf{x^{\prime}}$ is retained through rejection sampling to enhance diversity.:
    $\max_{\mathbf{x}\in S_{k}^{(t-1)}}\cos(\mathcal{M}_{\text{ebd}}(\mathbf{x^{\prime}}),\mathcal{M}_{\text{ebd}}(\mathbf{x}))<\tau,$ (7)
    where $\mathcal{M}_{\text{ebd}}(\cdot)$ indicates the output of the embedding layer of the given LLM, with $\tau$ as the similarity threshold. This process terminates when it reaches the predetermined iteration. The implementation of all parameters, along with their corresponding ablation studies, is provided in Appendix C.1.
The domain centroid is then computed from the final augmented set $\mathcal{S}_{k}$ using the model’s embedding, formalized as $\mathcal{C}_{k}=1/|\mathcal{S}_{k}|\cdot\sum_{\mathbf{x}_{i}\in\mathcal{S}_{k}}\mathcal{M}_{\text{ebd}}(\mathbf{x}_{i})$. This captures the LLM’s intrinsic feature space geometry while eliminating dependency on human annotations.
Domain-Aware Clustering
We then automatically construct pseudo-labels for the given data samples based on the previously synthesized centroids $\{C_{k}\}_{k=1}^{K}$. We perform constrained k-means clustering in the embedding space with $\arg\min_{\{S_{k}\}}\sum_{k=1}^{K}\sum_{\tilde{x}\in S_{k}}\|\mathcal{M}_{\text{ebd}}(\tilde{x})-\mathcal{C}_{k}\|^{2}$. This produces the seed dataset $\mathcal{D}_{\text{probe}}$ containing pseudo-labels $\{\tilde{y}_{i}\}_{i=1}$ where $\tilde{y}_{i}\in\{1,\ldots,K\}$, with model-induced and embedding-derived domain label assignments.
4.2 Training for Self-Rewarding Abilities
Entropy as a Diversity Proxy
While diversity metrics in Eqs. (2)-(3) directly quantify cross-sample distribution, we aim to design a reward probe that outputs with sample-level diversity level. Since models require sample-level processing that is unable to leverage pairwise relationships, direct training for diversity leads to poor performance in Appendix F.4. We instead use softmax confidence scores and predictive entropy as an implicit diversity proxy, reflecting model-aware data discriminability. This enables effective diversity approximation through a two-stage framework.
Stage 1: Domain Predictor
The proposed DaaR establishes model-aware domain discrimination abilities through a multi-layer perceptron (MLP) probe module, $\psi_{\text{dom}}$, attached to the hidden layer $\mathcal{M}(\tilde{x})$ of the LLMs. The probe will be trained meanwhile all the parameters of the LLM are frozen, achieving a preferable balance between effectiveness and cost, with detailed analysis regarding the choice of layers presented in Appendix G.2. Specifically, with pseudo-label $\tilde{y}$, we can compute domain predicted probabilities as:
$p_{k}(\tilde{x})=\text{softmax}\left(\psi_{\text{dom}}(\mathcal{M}(\tilde{x}))\right),\quad\psi_{\text{dom}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{K},$ (8)
$\psi_{\text{dom}}$ is optimized via cross-entropy loss $\mathcal{L}_{\text{dom}}=-\frac{1}{|\mathcal{D}_{\text{probe}}|}\sum_{(\tilde{x},\tilde{y})\in\mathcal{D}_{\text{probe}}}\sum_{k=1}^{K}\mathbb{I}_{[k=\tilde{y}]}\log p_{k}(\tilde{x})$, where $\mathbb{I}_{[k=\tilde{y}]}$ denotes the indicator function. We employ single-sample batches with the AdamW optimizer to prevent gradient averaging across domains. Training consistently converges and achieves 92.7% validation accuracy on domain classification, as shown in Fig. 2 (a).
Stage 2: Diversity Rewarding
Building on the stabilized domain discrimination probe, we quantify sample-level diversity through predictive entropy $H(\tilde{x})$. And to enable efficient reward computation during data selection, we then train another 5-layer MLP $\psi_{\text{div}}$ to directly estimate $H(\tilde{x})$ from $\mathcal{M}(\tilde{x})$:
$H(\tilde{x})=-\sum_{k=1}^{K}p_{k}(\tilde{x})\log p_{k}(\tilde{x}).\quad\hat{H}(\tilde{x})=\psi_{\text{div}}(\mathcal{M}(\tilde{x})),\psi_{\text{div}}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}.$ (9)
This diversity probe module $\psi_{\text{div}}$ shares $\psi_{\text{dom}}$’s architecture up to its final layer that replaced with the regression head, trained using entropy-scaled MSE loss $\mathcal{L}_{\text{div}}=\frac{1}{|\mathcal{D}_{\text{probe}}|}\sum_{\tilde{x}\in\mathcal{D}_{\text{probe}}}(\hat{H}(\tilde{x})-H(\tilde{x}))^{2}$. The module is also well-converged as shown in Fig. 2 (b).
Figure 2: Training loss and validation process of the two stages of DaaR on Qwen2-7B, more detailed results in Appendix H.3.
Data Selection: After training the module $\psi_{\text{div}}$, we can use its output to select data samples. Building on the theoretical insights in Sec. 3.4, data points that are closer to other centroids and more dispersed within their own centroid are more beneficial for enhancing the comprehensive capabilities of the model. Therefore, we use the predicted entropy score as a reward, selecting the top 20% with the highest scores as the final data subset for fine-tuning. The analysis of stability is shown in Appendix G.8.
5 Experiments
To validate the efficacy of DaaR, we conduct comprehensive experiments on data pools and benchmarks in Sec. 3.1, comparing SOTA baselines as follows with critical modifications: all domain-specific labels are deliberately stripped. This constraint mimics more challenging real-world scenarios and precludes direct comparison with data mixture methods requiring domain label prior.
Baselines
We use the following data selection methods for comprehensive evaluation: (1) Random Selection: traditional random sampling; (2) Instruction Len: measuring instruction complexity by token count  cao2023instruction ; (3) Alpagasus chen2024alpagasus : using ChatGPT for direct quality scoring of instruction pairs; (4-5) Instag lu2023instag : semantic analysis approach with Instag-C (complexity scoring via tag quantity) and Instag-D (diversity measurement through tag set expansion); (6) SuperFilter li2024super : response-loss-based complexity estimation using compact models; (7-9) Deita liu2023deita : model-driven evaluation with Deita-C (complexity scoring), Deita-Q (quality scoring), and Deita-D (diversity-aware selection). Detailed implementations for these baselines are in Appendix D.1.
Table 2: Evaluation results of Llama3.1-8B, Qwen2-7B, and Qwen2.5-7B across various downstream task benchmarks. DaaR demonstrates superiority in AVG compared to other baselines.
@@ 5 Experiments
Full (40K) 21.92 65.11 73.62 51.70 7.60 4.10 36.50 37.22
Random (8K) 21.99 64.83 74.72 55.70 14.50 5.10 24.09 37.27
Instruction Len 15.34 63.60 73.73 54.00 15.40 3.60 30.80 36.64
Alpagasus chen2024alpagasus 21.57 64.37 74.87 55.20 17.65 4.60 16.16 36.34
Instag-Best lu2023instag 18.12 64.96 74.01 55.70 15.50 4.80 37.81 38.70
SuperFilter li2024super 22.95 64.99 76.39 57.60 6.05 2.60 40.55 38.73
Deita-Best liu2023deita 15.58 64.97 74.21 55.00 13.05 4.60 34.46 37.41
DaaR (Ours) 20.08 64.55 74.88 54.80 15.30 4.70 37.50 38.83
@@
Full (40K) 15.61 58.75 72.51 73.80 31.30 51.70 67.38 53.01
Random (8K) 13.28 58.27 73.00 75.35 35.36 52.20 63.72 53.02
@@
Alpagasus chen2024alpagasus 13.67 57.94 73.04 73.90 32.30 51.40 63.41 52.24
Instag-Best lu2023instag 9.51 58.50 73.06 74.70 35.35 51.90 64.70 52.53
SuperFilter li2024super 19.16 58.98 72.99 73.70 30.10 52.40 58.85 52.31
Deita-Best liu2023deita 16.41 57.80 72.70 76.10 29.05 52.40 64.63 52.73
@@
Alpagasus chen2024alpagasus 10.40 57.87 72.92 77.20 18.75 61.80 65.55 52.07
Instag-Best lu2023instag 11.08 58.40 72.79 76.40 16.40 62.90 70.43 52.63
SuperFilter li2024super 13.54 58.51 72.89 79.30 11.35 39.50 65.25 48.62
Deita-Best liu2023deita 10.50 58.17 73.14 74.60 16.60 62.00 72.26 52.47